# Less is More: Skim Transformer for Light Field Image Super-resolution

### v2

@@ IV-D Model Efficiency
We assess the efficiency of SkimLFSR by benchmarking it against state-of-the-art methods at the 4×4$\times$4 × scale. To assess performance across a spectrum of model complexities, we vary the number of correlation blocks $N_{CB}$ from 4 to 20 in increments of two. The compared methods include the Transformer-based LFT [26], EPIT [24], LF-DET [25] and M2MT-Net [27], as well as the convolution-based DistgSSR [60], LFSSR-SAV [58] and HLFSR [61]. The results are presented in Table III and plotted in Fig. 6. In the plots, the average PSNR is used as the performance metric on the vertical axis. Three efficiency metrics are involved as the horizontal axis: (a) parameter number, (b) FLOPs, and (c) inference time. The parameter number is measured in millions (M), indicating model size, FLOPs (floating-point operations) are measured in billions (G), indicating theoretical computational complexity, and the inference time is measured in seconds, indicating the actual elapsed time of processing a LF image. Ideally, a method positioned closer to the top-left corner of each plot is considered more efficient, as it achieves higher performance with fewer computational resources.
@@
Our SkimLFSR network consistently excels across all three efficiency metrics. Notably, the most lightweight variant of SkimLFSR with $N_{CB}=4$ has already achieved a PSNR of 32.47 dB, outperforming all compared methods except M2MT-Net, and surpassing LF-DET by 0.06 dB. Remarkably, it does so with only 37% of the parameters, 35% of the FLOPs, and 28% of the inference time compared to LF-DET. This result aligns with our analysis in Section III-D and Table I, validating that compared to conventional spatial Transformers, Skim Transformers reduces computational redundancy through its skimming strategy while maintaining information integrity and emphasizing more pertinent information within the self-attention mechanism, embodying the “less is more” philosophy.
@@
Compared to the best existing method, M2MT-Net, the smallest SkimLFSR variant that surpasses it is with $N_{CB}=8$, achieving 32.80 dB, which is 0.09 dB higher. While the FLOPs and inference time are comparable (97% and 108%, respectively), this SkimLFSR model requires only 28% of the parameters, leading to a significantly lighter architecture. The reduced model size not only lowers the risk of overfitting but also enables easier scalability to deeper networks.
@@
Indeed, as $N_{CB}$ increases, SkimLFSR consistently improves in performance. The deepest variant with $N_{CB}=20$ reaches 33.06 dB, outperforming M2MT-Net and LF-DET by 0.35 dB and 0.65 dB, respectively. Remarkably, this model still uses only 67% of the parameters of M2MT-Net and is only 4% slower than LF-DET in inference time, while delivering significantly better results.
@@
These results demonstrate that SkimLFSR achieves a compelling balance between performance and efficiency. It is capable of delivering lightweight variants, such as those with $N_{CB}=4$ and $N_{CB}=8$, that offer strong performance under limited computational budgets. At the same time, it provides scalable, high-performance configurations such as $N_{CB}=20$, which significantly outperform existing methods while maintaining a manageable model size and runtime. This adaptability makes SkimLFSR a practical and powerful solution for a wide range of real-world and high-demand LFSR applications.
@@ IV-E In-depth Analysis
To better understand the underlying mechanism driving SkimLFSR’s superior performance, we further conduct a series of in-depth analyses on the challenging 4×4$\times$4 × scale.
@@ IV-E1 DSA Feature Visualization
$DSA_{1}$
$DSA_{2}$
$DSA_{1}-DSA_{2}$ $DSA_{2}-DSA_{1}$
Refer to caption Refer to caption Refer to caption Refer to caption
(a) Perforated_Metal_3
Refer to caption Refer to caption Refer to caption Refer to caption
(b) Building_Decoded
Refer to caption Refer to caption Refer to caption Refer to caption
(c) Lego Knights
Figure 7: Visualization of DSA feature maps. The first two columns display the outputs of the two DSA branches, respectively. The third and fourth columns present their differential activations, revealing the exclusive attention regions of each branch.
@@
Specifically, in Fig. 5 (a) Perforated_Metal_3, $DSA_{1}-DSA_{2}$ demonstrates strong activation on the perforated metal surface in the foreground, whereas $DSA_{2}-DSA_{1}$ predominantly highlights the background regions behind the metal board. This disparity separation is particularly noteworthy since the foreground and background regions exhibit markedly different disparity ranges, suggesting that SkimLFSR has developed an inherent capability to distinguish the perforated metal structure from the background.
The result in Fig. 5 (b) Building_Decoded reveals a different specialization pattern. Here, $DSA_{1}-DSA_{2}$ appears to attend preferentially to background regions, including building walls, sky regions, and grass fields. Conversely, $DSA_{2}-DSA_{1}$ exhibits a pronounced tendency to focus on structural boundaries, particularly edges of the building and its windows.
Interestingly, an opposite pattern is observed in Fig. 5 (c) Lego Knights. In this case, $DSA_{1}-DSA_{2}$ primarily attends to structural edges, such as the knights and studs, while $DSA_{2}-DSA_{1}$ focuses on background regions. This reversal can be attributed to the fundamental differences between samples Fig. 5 (b) and (c): the former originates from a Lytro Illum camera capturing a far scene, while the latter is synthetically generated with a completely different camera configuration and presents near objects in a simulated environment. As a result, these two DSA branches demonstrate adaptive behavior, attending to content with different characteristics as well as the specific camera configuration.
These observations validate our hypothesis that the multi-branch architecture effectively mitigates disparity entanglement issues, enabling SkimLFSR to develop adaptive disparity-aware capabilities that facilitate comprehensive scene understanding during reconstruction. It is worth noting that, despite being trained solely on the regression-based LFSR task, in this study, considering separation of the perforated surface in (a) and locating the structures of building in (b) and lego in (c), SkimLFSR demonstrates emergent behavior resembling classification, suggesting implicit learning of semantics akin to classification tasks, like depth estimation [5] and shape extraction [4], highlighting the effectiveness of the proposed disparity disentanglement mechanism.
@@
(a) EPIT (b) LF-DET (c) M2MT-Net (d) SkimLFSR (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption
CHI: 17.25 CHI: 6.48 CHI: 11.86 CHI: 41.27
Figure 8: Projected feature representation using t-SNE. Each plot contains 167 points, each of which represents a sample in the dataset. Colors indicate the light field camera configuration: Lytro Illum (blue), synthetic (green), and gantry-based (red). Higher CHI scores indicate better cluster separability.
@@
We further analyze SkimLFSR’s feature space using deep degradation representation (DDR), a technique originally proposed for interpreting single image super-resolution (SISR) models. In prior work [72, 73], it is discovered that by applying PCA [74] and t-SNE [75] to project the deep features into a two-dimensional space, well-trained SISR methods exhibit a strong clustering trend based on degradation types, suggesting that these methods have implicitly learned degradation as semantics, even though such information is not explicitly provided during training.
@@
Inspired by this observation, we hypothesize that an effective LFSR method should similarly exhibit patterns related to disparity information in LF images. To validate this hypothesis, we perform the same technique to the three LFSR methods: EPIT [24], LF-DET [25], and M2MT-Net [27], on all the 167 samples from the BasicLFSR framework [60], as shown in Fig. 8. These samples are categorized into three camera configurations: the Lytro Illum camera group (blue), EPFL [68], INRIA [71]; the synthetic group (green), HCInew [69], HCIold [70]; and the gantry-based camera group (red), STFgantry [21]. Each group is characterized by distinct disparity distributions due to different imaging setups: the Lytro Illum and gantry-based cameras employ different microlens configurations, while synthetic datasets simulate yet another unique angular distribution. To quantitatively assess the clustering trend, we compute the Calinski-Harabasz Index (CHI) [76] for each method as the metric, shown below the respective plots. The CHI score measures the ratio of between-cluster heterogeneity to within-cluster homogeneity; a higher CHI score indicates better cluster separability in the projected feature space.
@@
Despite the clear disparity distinctions across camera configuration groups, the feature visualizations in Fig. 8 show that the compared methods exhibit no evident clustering trend, suggesting a lack of disparity-aware representation. Among them, (b) LF-DET demonstrates the weakest separation (CHI: 6.48), with all three groups dispersed across the feature space. (a) EPIT (CHI: 17.25) and (c) M2MT-Net (CHI: 11.86) display a slight clustering trend by camera configurations, but substantial overlaps persist, with numerous outliers from each group intermingled with others. These results point to the potential disparity entanglement issue in these methods, where disparity-related cues are not effectively disentangled or captured in the learned feature representations.
In contrast, (d) SkimLFSR exhibits a clear and well-defined clustering pattern, with distinct boundaries separating the three groups and only minor overlaps. Its CHI score of 41.27 is substantially higher than those of the other methods, indicating that SkimLFSR has effectively learned to organize features based on disparity information.
This outcome strongly echoes our earlier observations in Fig. 7, where SkimLFSR demonstrates classification-like behavior despite being trained solely on a regression-based LFSR task. Whereas the previous analysis highlights SkimLFSR’s ability to distinguish scene depths within a single LF image, the current t-SNE study reveals its capacity to differentiate camera configurations across LF images. Since scene depth and camera configuration are the two primary factors influencing disparity variations in LF data, these results collectively indicate that SkimLFSR has developed a latent discriminative capability toward disparity information, i.e., disparity-aware, which finally contributes to its superior performance in LFSR tasks.
IV-E3 Ablation study of Skimmed SAI Sets
@@
TABLE IV: Comparison of different skimmed SAI sets for SkimLFSR ($N_{CB}=6$) of 4×4$\times$4 × LFSR.
Variant (a) (b) (c) (d) (e) (f) (g) (h) (i)
@@
The configuration of the skimmed SAI sets in the DSA branches serves as critical prior knowledge for targeted disparity modeling in the Skim Transformer. To evaluate its impact, we conduct an ablation study, with results presented in Table IV.
We adopt the $N_{CB}=6$ model as the baseline, representing the full version of Skim Transformers in Column (a).
@@
Columns (b) and (c) isolate the two individual skimmed SAI sets (blue and red) used in the baseline. Each subset underperforms the combined approach by 0.11 dB and 0.13 dB, respectively, demonstrating the advantage of combining different SAI sets to capture a broader disparity range.
@@
Columns (d), (e), and (f) explore variants of (a), (b), and (c) using only half the number of SAIs. These configurations, which sample along a single diagonal line with a biased disparity coverage, lead to further degradation in performance: a drop of 0.18 dB for the dual-branch variant (Column (a) vs. (d)), and 0.28 dB for the single-branch variants (Column (a) vs. (e) and (f)). These results highlight the importance of both the quantity and spatial distribution of SAIs, as they directly influence the disparity coverage and, consequently, the model’s reconstruction quality.

### v3

@@ IV-D Model Efficiency
We assess the efficiency of SkimLFSR by benchmarking it against state-of-the-art methods at the $4\times$ scale. To assess performance across a spectrum of model complexities, we vary the number of correlation blocks $N_{CB}$ from 4 to 20 in increments of two. The compared methods include the Transformer-based LFT, EPIT, LF-DET, and M2MT-Net, as well as the convolution-based DistgSSR, LFSSR-SAV, and HLFSR. The results are plotted in Fig. 6. In the plots, the average PSNR is used as the performance metric on the vertical axis. Three efficiency metrics are involved as the horizontal axis: (a) parameter number, (b) FLOPs, and (c) inference time. Ideally, a method positioned closer to the top-left corner of each plot is considered more efficient, as it achieves higher performance with fewer computational resources. The detailed numerical results are provided in the supplementary material.
@@
Our SkimLFSR network consistently excels across all three efficiency metrics. Notably, the most lightweight variant of SkimLFSR with $N_{CB}=4$ has already achieved a PSNR of 32.47 dB, outperforming all compared methods except M2MT-Net, and surpassing LF-DET by 0.06 dB. Remarkably, it does so with only 37% of the parameters, 35% of the FLOPs, and 28% of the inference time compared to LF-DET. This result aligns with our analysis in Section III-D and Table I, validating that through the skimming strategy, Skim Transformers reduces computational redundancy while maintaining information integrity and emphasizing more pertinent information, embodying the “less is more” philosophy.
@@
Compared to the best existing method, M2MT-Net, the smallest SkimLFSR variant that surpasses it is with $N_{CB}=8$, achieving 32.80 dB, which is 0.09 dB higher. While the FLOPs and inference time are comparable (97% and 108%, respectively), this SkimLFSR model requires only 28% of the parameters, leading to a significantly lighter architecture. The reduced model size not only lowers the risk of overfitting but also enables easier scalability to deeper networks. Meanwhile, SkimLFSR also surpasses the Mamba-based LFMamba, and it requires only 49% of the parameters and 45% of the FLOPs, and is 19% faster in inference time.
@@
Indeed, as $N_{CB}$ increases, SkimLFSR consistently improves in performance. The variant with $N_{CB}=20$ reaches 33.06 dB, outperforming M2MT-Net and LF-DET by 0.35 dB and 0.65 dB, respectively. Remarkably, this model still uses only 67% of the parameters of M2MT-Net, and is only 4% slower than LF-DET and 19% faster than LFMamba in inference time, while delivering significantly better results. After this point, the model performance starts to drop when adding more correlation blocks ($N_{CB}=22$ and $24$), indicating that the model has reached capacity saturation.
@@
In summary, these results demonstrate that SkimLFSR achieves a compelling balance between performance and efficiency. It is capable of delivering lightweight variants, such as $N_{CB}=4$ and $N_{CB}=8$, that offer strong performance under limited computational budgets. At the same time, it provides scalable configurations such as $N_{CB}=20$, which significantly outperform existing methods while maintaining a manageable model size and runtime. This adaptability makes SkimLFSR a practical and powerful solution for a wide range of real-world and high-demand LFSR applications.
Refer to caption $DSA_{1}$ Refer to caption Refer to caption Refer to caption
Refer to caption $DSA_{2}$ Refer to caption Refer to caption Refer to caption
$DSA_{1}-DSA_{2}$ Refer to caption Refer to caption Refer to caption
$DSA_{2}-DSA_{1}$ Refer to caption Refer to caption Refer to caption
(a) Perforated_Metal_3 (b) Building_Decoded (c) Lego Knights
Figure 7: Visualization of DSA feature maps. The first two columns display the two DSA branches’ output. The third and fourth columns present their differential activations, revealing the exclusive attention of each branch.
(a) EPIT (b) LF-DET (c) M2MT-Net (d) SkimLFSR (Ours)
CHI: 17.25 CHI: 6.48 CHI: 11.86 CHI: 41.27
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Projected feature representation using t-SNE. Each plot contains 167 points, each of which represents a sample in the dataset.
@@ IV-E In-depth Analysis
To better understand the underlying mechanism driving SkimLFSR’s superior performance, we further conduct a series of in-depth analyses on the challenging $4\times$ scale.
@@ IV-E1 DSA Feature Visualization
Specifically, in (a) Perforated_Metal_3, $DSA_{1}-DSA_{2}$ demonstrates strong activation on the perforated metal surface in the foreground, whereas $DSA_{2}-DSA_{1}$ predominantly highlights the background regions behind the metal board. This disparity separation is particularly noteworthy since the foreground and background regions exhibit markedly different disparity ranges, suggesting that SkimLFSR has developed an inherent capability to distinguish the perforated metal structure from the background.
The result in (b) Building_Decoded reveals a different specialization pattern. Here, $DSA_{1}-DSA_{2}$ appears to attend preferentially to background regions, including building walls, sky regions, and grass fields. Conversely, $DSA_{2}-DSA_{1}$ exhibits a pronounced tendency to focus on structural boundaries, particularly edges of the building and its windows.
Interestingly, an opposite pattern is observed in (c) Lego Knights. In this case, $DSA_{1}-DSA_{2}$ primarily attends to structural edges, such as the knights and studs, while $DSA_{2}-DSA_{1}$ focuses on background regions. This reversal can be attributed to the fundamental differences between samples (b) and (c): the former originates from a Lytro Illum camera capturing a far scene, while the latter is synthetically generated with a completely different camera configuration and presents near objects in a simulated environment. As a result, these two DSA branches demonstrate adaptive behavior, attending to content with different characteristics as well as the specific camera configuration.
Notably, despite being trained solely on the regression-based LFSR task, considering separation of the perforated surface in (a) Perforated_Metal_3 and locating the structures of building in (b) Building_Decoded and lego in (c) Lego Knights, SkimLFSR demonstrates emergent behavior resembling classification, suggesting implicit learning of semantics akin to classification tasks, like depth estimation [43] and shape extraction [18], highlighting that SkimLFSR has developed adaptive disparity-aware capabilities that facilitate comprehensive scene understanding during reconstruction.
@@
We further analyze SkimLFSR’s feature space using deep degradation representation (DDR), a technique originally proposed for interpreting single image super-resolution (SISR) models. In prior work [32, 24], it is discovered that by applying PCA [20] and t-SNE [33] to project the deep features into a two-dimensional space, well-trained SISR methods exhibit a clustering trend based on degradation types, suggesting that these methods have implicitly learned degradation as semantics, even though such information is not explicitly provided during training.
@@
Inspired by this observation, we hypothesize that an effective LFSR method should similarly exhibit patterns related to disparities. To validate it, we perform the same technique on the three LFSR methods: EPIT [28], LF-DET [9], and M2MT-Net [21], on all the 167 samples, as shown in Fig. 8. These samples are categorized into three camera configurations: the Lytro Illum camera group (blue), EPFL [37], INRIA [25]; the synthetic group (green), HCInew [19], HCIold [48]; and the gantry-based camera group (red), STFgantry [39]. Each group is characterized by distinct disparity distributions due to different imaging setups: the Lytro Illum and gantry-based cameras employ different microlens configurations, while synthetic datasets simulate yet another unique angular distribution. The Calinski-Harabasz Index (CHI) [3] is computed for each method to quantify the clustering trend, with higher scores indicating better separability of the clusters.
@@
Despite the clear disparity distinctions across camera configuration groups, the compared methods exhibit no evident clustering trend, suggesting a lack of disparity-aware representation. LF-DET demonstrates the weakest separation (CHI: 6.48), with all three groups dispersed across the feature space. EPIT (CHI: 17.25) and M2MT-Net (CHI: 11.86) display a slight clustering trend, but substantial overlaps persist, with numerous outliers from each group intermingled with others. These results point to the potential disparity entanglement issue in the learned feature representations of these methods.
@@
TABLE III: Comparison of different skimmed SAI sets for SkimLFSR ($N_{CB}=6$) in $4\times$ LFSR.
Variants (a) (b) (c) (d) (e) (f) (g) (h) (i)
@@
Figure 9: Comparison of performance of the double-branch baseline and the single-branch variant (Variant (a) and Variant (i) in Table III) varying the number of correlation blocks $N_{CB}$.
In contrast, SkimLFSR exhibits a clear and well-defined clustering pattern, with distinct boundaries separating the three groups and only minor overlaps. Its CHI score of 41.27 is substantially higher than those of the other methods, indicating that SkimLFSR has effectively learned to organize features based on disparity information.
This outcome strongly echoes our earlier observations in Fig. 7, where SkimLFSR demonstrates classification-like behavior despite being trained solely on a regression-based LFSR task. Whereas the previous analysis highlights SkimLFSR’s ability to distinguish scene depths within a single LF image, the current t-SNE study reveals its capacity to differentiate camera configurations across LF images. Since scene depth and camera configuration are the two primary factors influencing disparity variations, these studies collectively prove that SkimLFSR has developed a latent discriminative capability toward disparity information, i.e., disparity-aware, which finally contributes to its superior performance in LFSR.
IV-E3 Ablation study on Skimmed SAI Sets
@@
The configuration of the skimmed SAI sets in the DSA branches serves as critical prior knowledge for targeted disparity modeling in the Skim Transformer. To evaluate its impact, we conduct an ablation study, with results presented in Table III. We adopt the $N_{CB}=6$ model as the baseline, representing the full version of Skim Transformers in Variant (a).
@@
Variants (b) and (c) isolate the two individual skimmed SAI sets (blue and red) used in the baseline. Each subset underperforms the combined variant by 0.11 dB and 0.13 dB, respectively, demonstrating the advantage of combining different SAI sets to capture a broader disparity range.
@@
Variants (d), (e), and (f) explore variants of Variant (a), (b), and (c) using only half the number of SAIs. These configurations, which sample along a single diagonal line with a biased disparity coverage, lead to further degradation in performance: a drop of 0.18 dB for the dual-branch variant (Variant (a) vs. (d)), and 0.28 dB for the single-branch variants (Variant (a) vs. (e) and (f)). These results highlight the importance of both the quantity and spatial distribution of SAIs, as they directly influence the disparity coverage and, consequently, the model’s reconstruction quality.
TABLE IV: Comparison of PSNR/SSIM on $7\times 7$ SAIs and $4\times$ LFSR. In the first group, the best and second-best results are highlighted in bold and underlined, respectively. In the second and third groups, the difference between the SkimLFSR baseline model and its variants is displayed below PSNR/SSIM.
Methods (#Train/#Test samples) EPFL (70/10) HCInew (20/4) HCIold (10/2) INRIA (35/5) STFgantry (9/2) Average
1. Qualitative Comparison (Trained on $7\times 7$ SAIs)
LF-DET [9] 30.07/0.9314 31.94/0.9286 38.10/0.9756 31.93/0.9578 32.57/0.9604 32.92/0.9508
M2MT-Net [21] 30.14/0.9314 31.80/0.9280 38.22/0.9761 31.99/0.9579 32.68/0.9606 32.97/0.9508
SkimLFSR (Ours) [Uncaptioned image] 30.40/0.9365 32.06/0.9304 38.33/0.9765 32.11/0.9601 32.96/0.9631 33.15/0.9532
+0.26/+0.0051 +0.12/+0.0018 +0.11/+0.0004 +0.12/+0.0022 +0.28/+0.0025 +0.18/+0.0024
2. Ablation Study on Skimmed SAI Sets (Trained on $7\times 7$ SAIs)
(a) [Uncaptioned image] 30.30/0.9357 32.04/0.9300 38.26/0.9763 31.96/0.9596 32.81/0.9625 33.07/0.9528
-0.10/-0.0008 -0.02/-0.0004 -0.07/-0.0002 -0.15/-0.0005 -0.15/-0.0006 -0.08/-0.0004
(b) [Uncaptioned image] 30.26/0.9352 32.03/0.9299 38.25/0.9762 31.94/0.9595 32.85/0.9628 33.07/0.9528
-0.14/-0.0013 -0.03/-0.0005 -0.08/-0.0003 -0.17/-0.0006 -0.11/-0.0003 -0.08/-0.0004
(c) [Uncaptioned image] 30.22/0.9354 32.02/0.9298 38.24/0.9763 31.92/0.9594 32.81/0.9626 33.04/0.9527
-0.18/-0.0011 -0.04/-0.0006 -0.09/-0.0002 -0.19/-0.0007 -0.15/-0.0005 -0.11/-0.0005
3. Cross-angular-resolution Generalization (Trained on $5\times 5$ SAIs $\rightarrow$ and tested on $7\times 7$ SAIs)
DistgSSR [46] N/A N/A N/A N/A N/A N/A
HLFSR [40] N/A N/A N/A N/A N/A N/A
LFT [29] 28.67/0.9112 30.15/0.8989 36.22/0.9636 30.66/0.9447 29.54/0.9220 31.05/0.9281
EPIT [28] 29.58/0.9230 31.63/0.9250 37.93/0.9749 31.49/0.9538 32.40/0.9585 32.61/0.9470
LF-DET [9] 29.68/0.9250 31.58/0.9239 37.90/0.9745 31.51/0.9541 32.14/0.9567 32.56/0.9469
LFMamba [51] 29.76/0.9270 31.34/0.9237 37.18/0.9717 31.63/0.9545 31.55/0.9535 32.29/0.9461
M2MTNet [21] N/A N/A N/A N/A N/A N/A
(d) [Uncaptioned image] $\rightarrow$ [Uncaptioned image] 30.21/0.9328 31.96/0.9284 38.18/0.9757 31.99/0.9581 32.71/0.9610 33.01/0.9512
-0.19/-0.0037 -0.10/-0.0020 -0.15/-0.0008 -0.12/-0.0020 -0.25/-0.0021 -0.14/-0.0020
(e) [Uncaptioned image] $\rightarrow$ [Uncaptioned image] 30.32/0.9337 31.98/0.9284 38.26/0.9762 32.07/0.9584 32.73/0.9610 33.07/0.9515
-0.08/-0.0028 -0.08/-0.0020 -0.07/-0.0003 -0.04/-0.0017 -0.23/-0.0021 -0.08/-0.0017
(f) [Uncaptioned image] $\rightarrow$ [Uncaptioned image] 30.35/0.9338 31.95/0.9273 38.21/0.9758 32.05/0.9582 32.66/0.9602 33.05/0.9511
-0.05/-0.0027 -0.11/-0.0031 -0.12/-0.0007 -0.06/-0.0019 -0.30/-0.0029 -0.10/-0.0021